# Nyström Kernel Stein Discrepancy

### v2

@@ Appendix C Derivation of (9)
$\sup_{\left\|{\mathbf{f}}\right\|_{\mathcal{H}_{p,m}}\leq 1}\left\langle{\mathbf{f},\mathbb{E}_{X\sim\hat{\mathbb{Q}}_{n}}\bm{\xi}_{p}(X)}\right\rangle_{\mathcal{H}_{k}^{d}}\stackrel{{\scriptstyle(i)}}{{=}}\sup_{\left\|{\mathbf{f}}\right\|_{\mathcal{H}_{p,m}}^{2}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\left\langle{\mathbf{f},\bm{\xi}_{p}(\mathbf{x}_{i})}\right\rangle_{\mathcal{H}_{k}^{d}}$
=$\stackrel{{\scriptstyle(ii)}}{{=}}\sup_{\bm{\alpha}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m}\bm{\alpha}\leq 1}\frac{1}{n}\sum_{i=1}^{n}\left\langle{\sum_{j=1}^{m}\alpha_{i}\bm{\xi}_{p}(\tilde{\mathbf{x}}_{j}),\bm{\xi}_{p}(\mathbf{x}_{i})}\right\rangle_{\mathcal{H}_{k}^{d}}$
=$\stackrel{{\scriptstyle(iii)}}{{=}}\sup_{\bm{\alpha}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m}\bm{\alpha}\leq 1}\sum_{j=1}^{m}\alpha_{i}\underbrace{\frac{1}{n}\sum_{i=1}^{n}h_{p}(\tilde{\mathbf{x}}_{j},\mathbf{x}_{i})}_{=:\beta_{p,j}}\stackrel{{\scriptstyle(iv)}}{{=}}\sup_{\bm{\alpha}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m}\bm{\alpha}\leq 1}\bm{\alpha}^{\mkern-1.5mu\mathsf{T}}\bm{\beta}_{p}\stackrel{{\scriptstyle(v)}}{{=}}\sup_{\left\|{\bm{\gamma}}\right\|_{2}^{2}\leq 1}\bm{\gamma}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-1/2}\bm{\beta}_{p}$
=$\stackrel{{\scriptstyle(vi)}}{{=}}\sqrt{\bm{\beta}_{p}^{\mkern-1.5mu\mathsf{T}}\mathbf{K}_{h_{p},m,m}^{-1}\bm{\beta}_{p}},$
where $(i)$ follows from the definition of the sample distribution ^$\hat{\mathbb{Q}}_{n}$ and linearity of the inner product, $(ii)$ uses that $\mathbf{f}\in\mathcal{H}_{p,m}$, and $(iii)$ follows from the linearity of the inner product and the definition of $h_{p}$. By defining $\bm{\beta}_{p}=(\beta_{p,j})_{j=1}^{m}=\frac{1}{n}\mathbf{K}_{h_{p},m,n}\bm{1}_{n}$, we obtain $(iv)$; in $(v)$, we introduce $\bm{\gamma}=\mathbf{K}_{h_{p},m,m}^{1/2}\bm{\alpha}$, and $(vi)$ is implied by the Cauchy-Schwarz inequality.
Appendix D External statements
@@
This section collects the external statements that we use. Lemma D.1 gives equivalent norms for $f\otimes f$. We collect properties of Orlicz norms in Lemma D.2. Theorem D.3 is about the concentration of the empirical covariance, and Theorem D.4 recalls Bernstein’s inequality for separable Hilbert spaces. Theorem D.5 is a concentration result for bounded random variables in a separable Hilbert space. Theorem D.6 recalls the combination of Rudi et al. [2015, Proposition 3 and Proposition 7].
Lemma D.1 (Lemma B.8; Sriperumbudur and Sterge 2022).
@@
We refer to the following sources for the items in Lemma D.2. Item 1 is Vershynin [2018, Lemma 2.6.8], Item 2 is Vershynin [2018, Exercise 2.7.10], Item 3 recalls van der Vaart and Wellner [1996, p. 95], and Item 4 is Vershynin [2018, Lemma 2.7.6].
Lemma D.2 (Collection of Orlicz properties).
@@
    $\left\|{X-\mathbb{E}X}\right\|_{\psi_{2}}\leq\left\|{X}\right\|_{\psi_{2}}+\left\|{\mathbb{E}X}\right\|_{\psi_{2}}\lesssim\left\|{X}\right\|_{\psi_{2}}.$ (D.55)
@@
    $\left\|{X-\mathbb{E}X}\right\|_{\psi_{1}}\leq\left\|{X}\right\|_{\psi_{1}}+\left\|{\mathbb{E}X}\right\|_{\psi_{1}}\lesssim\left\|{X}\right\|_{\psi_{1}}.$ (D.56)
@@
    $\left\|{X^{2}}\right\|_{\psi_{1}}=\left\|{X}\right\|_{\psi_{2}}^{2}.$ (D.57)
Theorem D.3 (Theorem 9; Koltchinskii and Lounici 2017).
@@
$\left\|{\hat{C}_{n}-C}\right\|_{\mathrm{op}}\leq c\left\|{C}\right\|_{\mathrm{op}}\max\left(\sqrt{\frac{r(C)}{n}},\;\sqrt{\frac{\log(1/\delta)}{n}}\right),$ (D.58)
@@
Theorem D.4 (Bernstein bound for separable Hilbert spaces; Theorem 3.3.4; Yurinsky 1995).
@@
Theorem D.5 (Concentration in separable Hilbert spaces; Lemma E.1; Chatalic et al. 2022).
@@
The following combination of Rudi et al. [2015, Proposition 3 and Proposition 7] is standard. In the following, we adapt it to our notation and give a short proof to ensure self-containedness.
Theorem D.6 (Proposition 3, Proposition 7; Rudi et al. 2015).
Let $C_{\mathbb{Q},h_{p}}=\int_{\mathbb{R}^{d}}h_{p}\left(\cdot,{\mathbf{x}}\right)\otimes h_{p}\left(\cdot,{\mathbf{x}}\right)\mathrm{d}\mathbb{Q}(\mathbf{x})$, $C_{\mathbb{Q},h_{p},\lambda}=C_{\mathbb{Q},h_{p}}+\lambda I$ ($\lambda>0$), $\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}=\frac{1}{m}\sum_{i=1}^{m}h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\otimes h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)$. For $\lambda>0$, let $\beta(\lambda)=\lambda_{\text{max}}\left(C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\left(C_{\mathbb{Q},h_{p}}-\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{-1/2}\right)$, $\left(\tilde{\mathbf{x}}_{i}\right)_{i=1}^{m}$ are the Nyström points, as defined in Section 4.1, $\mathcal{H}_{h_{p},m}:=\mathrm{span}\left\{h_{p}\left(\cdot,{\tilde{\mathbf{x}}_{i}}\right)\mid i\in[m]\right\}$, and $P_{\mathcal{H}_{h_{p},m}}$ denotes the orthogonal projection onto $\mathcal{H}_{h_{p},m}$. Then, it holds that
$\left\|{\left(I-P_{\mathcal{H}_{h_{p},m}}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)},$ (D.59)
when $\beta(\lambda)<1$.
Proof.
$\left\|{\left(I-P_{h_{p},m}\right)C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(a)}}{{\leq}}\lambda\left\|{\left(Z^{*}_{m}Z_{m}+\lambda I\right)^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\stackrel{{\scriptstyle(b)}}{{=}}\lambda\left\|{\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}$ (D.60)
where (a) follows from Rudi et al. [2015, Proposition 3] with $X:=C_{\mathbb{Q},h_{p},\lambda}^{1/2}$ therein. (b) is by Sterge and Sriperumbudur [2022, Lemma A.7(iv)].
Applying the second inequality in the statement of Rudi et al. [2015, Proposition 7] to (D.60) (we specialize $A:=\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p}}$ and $B:=C_{\mathbb{Q},h_{p}}$ therein), we obtain
$\lambda\left\|{\tilde{C}_{\hat{\mathbb{Q}}_{n},h_{p},\lambda}^{-1/2}C_{\mathbb{Q},h_{p},\lambda}^{1/2}}\right\|_{\mathrm{op}}^{2}\leq\frac{\lambda}{1-\beta(\lambda)},$ (D.61)
when $\beta(\lambda)<1$. The combination of (D.60) and (D.61) yields the stated claim. ∎

### v4

@@ Appendix B AUXILIARY RESULTS
    For any $x\in\mathcal{H}$, one has $\left\langle{x,X^{*}AXx}\right\rangle_{\mathcal{H}}=\left\langle{Xx,AXx}\right\rangle_{\mathcal{H}}\stackrel{{\scriptstyle(\dagger)}}{{\leq}}\left\langle{Xx,BXx}\right\rangle_{\mathcal{H}}=\left\langle{x,X^{*}BXx}\right\rangle_{\mathcal{H}}$; ()($\dagger$)( † ) follows from $A\preccurlyeq B$ applied to $Xx$.
  2. 2.
    We apply (1.) with $X=B^{-1/2}$.
  3. 3.
    We have $B^{-1}=B^{-1/2}IB^{-1/2}\succcurlyeq B^{-1/2}BB^{-1/2}=I$, where we used (1.) in the second step.
  4. 4.
    By (2.), it holds that $B^{-1/2}AB^{-1/2}\preccurlyeq I$, from which (3.) implies that $B^{1/2}A^{-1}B^{1/2}\succcurlyeq I$. Now apply (1.) with $X=B^{-1/2}$ to obtain the stated result.
  5. 5.
    The $C^{*}$-property, the definition of the adjoint and that of the operator norm yield
    $\left\|{CX}\right\|_{\mathrm{op}}^{2}=\left\|{X^{*}C^{*}CX}\right\|_{\mathrm{op}}=\sup_{\left\|{x}\right\|_{\mathcal{H}}=1}\left\langle{x,X^{*}C^{*}CXx}\right\rangle_{\mathcal{H}}=\sup_{\left\|{x}\right\|_{\mathcal{H}}=1}\left\langle{Xx,C^{*}CXx}\right\rangle_{\mathcal{H}}$ (B.75)
    $\leq\sup_{\left\|{x}\right\|_{\mathcal{H}}=1}\left\langle{Xx,D^{*}DXx}\right\rangle_{\mathcal{H}}=\sup_{\left\|{x}\right\|_{\mathcal{H}}=1}\left\langle{x,X^{*}D^{*}DXx}\right\rangle_{\mathcal{H}}=\left\|{X^{*}D^{*}DX}\right\|_{\mathrm{op}}=\left\|{DX}\right\|_{\mathrm{op}}^{2},$ (B.76)
    which, after taking the positive square root, proves the claim. ∎
The following lemma collects some inequalities for the trace and operator norms of covariance operators. Many of these are known and frequently employed without proof; we provide proofs here for completeness.
Lemma B.7 (Covariance operator inequalities).
Let $\mathcal{H}$ be a separable Hilbert space, $X\sim\mathbb{Q}\in\mathcal{M}_{1}^{+}(\mathcal{H})$, $C_{\mathbb{Q}}=\mathbb{E}\left[X\otimes X\right]$, $C_{\mathbb{Q},\lambda}=C_{\mathbb{Q}}+\lambda I$, and let $r(\cdot)=\frac{\operatorname{tr}\left(\cdot\right)}{\left\|{\cdot}\right\|_{\mathrm{op}}}$ be defined on trace-class operators. Assume that $0<\lambda\leq\left\|{C_{\mathbb{Q}}}\right\|_{\mathrm{op}}$. Then, the following hold.
  1. 1.
    $\frac{1}{2}r\left(C_{\mathbb{Q}}\right)\leq\operatorname{tr}\left(C_{\mathbb{Q},\lambda}^{-1}C_{\mathbb{Q}}\right)\leq\frac{\operatorname{tr}\left(C_{\mathbb{Q}}\right)}{\lambda}$,
  2. 2.
    $\frac{1}{2}\leq\left\|{C_{\mathbb{Q},\lambda}^{-1/2}C_{\mathbb{Q}}C_{\mathbb{Q},\lambda}^{-1/2}}\right\|_{\mathrm{op}}<1$, and
  3. 3.
    $r\left(C_{\mathbb{Q},\lambda}^{-1/2}C_{\mathbb{Q}}C_{\mathbb{Q},\lambda}^{-1/2}\right)\leq\frac{2\operatorname{tr}\left(C_{\mathbb{Q}}\right)}{\lambda}$.
Proof.
Let $\left(\lambda_{i}\right)_{i\in I}$ denote the eigenvalues of $C_{\mathbb{Q}}$, with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq 0$.
  1. 1.
    The first inequality follows from $\operatorname{tr}\left(C_{\mathbb{Q},\lambda}^{-1}C_{\mathbb{Q}}\right)=\sum_{i\in I}\frac{\lambda_{i}}{\lambda_{i}+\lambda}\geq\sum_{i\in I}\frac{\lambda_{i}}{2\left\|{C_{\mathbb{Q}}}\right\|_{\mathrm{op}}}=\frac{\operatorname{tr}\left(C_{\mathbb{Q}}\right)}{2\left\|{C_{\mathbb{Q}}}\right\|_{\mathrm{op}}}$. The second one is footnote 3.
  2. 2.
    For the first inequality, observe that $\left\|{C_{\mathbb{Q},\lambda}^{-1/2}C_{\mathbb{Q}}C_{\mathbb{Q},\lambda}^{-1/2}}\right\|_{\mathrm{op}}=\frac{\lambda_{1}}{\lambda_{1}+\lambda}\stackrel{{\scriptstyle(\dagger)}}{{\geq}}\frac{1}{2}$, where $(\dagger)\Leftrightarrow 2\lambda_{1}\geq\lambda_{1}+\lambda\Leftrightarrow{(\left\|{C_{\mathbb{Q}}}\right\|_{\mathrm{op}}=)}\lambda_{1}\geq\lambda$, which holds by assumption. The second one is implied as $\frac{\lambda_{1}}{\lambda_{1}+\lambda}\stackrel{{\scriptstyle(\dagger)}}{{<}}1$, where $(\dagger)\Leftrightarrow\lambda_{1}<\lambda_{1}+\lambda\Leftrightarrow 0<\lambda$; this condition was again assumed.
  3. 3.
    We upper bound the numerator of $r(C_{\mathbb{Q},\lambda}^{-1/2}C_{\mathbb{Q}}C_{\mathbb{Q},\lambda}^{-1/2})$ by (1.) after, using the cyclic invariance of the trace, we rewrite it as $\operatorname{tr}\left(C_{\mathbb{Q},\lambda}^{-1/2}C_{\mathbb{Q}}C_{\mathbb{Q},\lambda}^{-1/2}\right)=\operatorname{tr}\left(C_{\mathbb{Q},\lambda}^{-1}C_{\mathbb{Q}}\right)$, and lower bound the denominator by (2.).∎
Appendix C EXTERNAL STATEMENTS
@@
This section collects the external statements that we use. Lemma C.1 gives equivalent norms for $f\otimes f$. We collect properties of Orlicz norms in Lemma C.2. Theorem C.3 is about the concentration of the empirical covariance, and Theorem C.4 recalls Bernstein’s inequality for separable Hilbert spaces. Theorem C.5 is a concentration result for bounded random variables in a separable Hilbert space.
Lemma C.1 (Lemma B.8; Sriperumbudur and Sterge 2022).
@@
We refer to the following sources for the items in Lemma C.2. Item 1 is Vershynin [2018, Lemma 2.6.8], Item 2 is Vershynin [2018, Exercise 2.7.10], Item 3 recalls van der Vaart and Wellner [1996, p. 95], and Item 4 is Vershynin [2018, Lemma 2.7.6].
Lemma C.2 (Collection of Orlicz properties).
@@
    $\left\|{X-\mathbb{E}X}\right\|_{\psi_{2}}\leq\left\|{X}\right\|_{\psi_{2}}+\left\|{\mathbb{E}X}\right\|_{\psi_{2}}\lesssim\left\|{X}\right\|_{\psi_{2}}.$ (C.77)
@@
    $\left\|{X-\mathbb{E}X}\right\|_{\psi_{1}}\leq\left\|{X}\right\|_{\psi_{1}}+\left\|{\mathbb{E}X}\right\|_{\psi_{1}}\lesssim\left\|{X}\right\|_{\psi_{1}}.$ (C.78)
@@
    $\left\|{X^{2}}\right\|_{\psi_{1}}=\left\|{X}\right\|_{\psi_{2}}^{2}.$ (C.79)
Theorem C.3 (Theorem 9; Koltchinskii and Lounici 2017).
@@
$\left\|{\hat{C}_{n}-C}\right\|_{\mathrm{op}}\leq c\left\|{C}\right\|_{\mathrm{op}}\max\left(\sqrt{\frac{r(C)}{n}},\;\sqrt{\frac{\log(1/\delta)}{n}}\right),$ (C.80)
@@
Theorem C.4 (Bernstein bound for separable Hilbert spaces; Theorem 3.3.4; Yurinsky 1995).
@@
Theorem C.5 (Concentration in separable Hilbert spaces; Lemma E.1; Chatalic et al. 2022).
@@
Appendix D ADDITIONAL EXPERIMENTS
In this section, we collect additional numerical results. Section D.1 discusses the trade-off between power and runtime of the tested approaches. Section D.2 shows the impact of the size of the Nyström sample.
D.1 Runtime vs. Power
Based on the experimental setup in Section 5, we performed an additional set of experiments to contrast runtime and power. We repeated each setup for $100$ rounds to obtain the given power and average runtime. The quadratic-time approaches are considered as baseline.
Figure 2: Runtime and power trade-off of the tested approximations.
Laplace vs. standard normal.
We fix $d=15$, $m=4\sqrt{n}$, and vary $n\in\{500,1000,2000\}$. The remaining parameters match the ones stated in Section 5.
Figure 2(a) summarizes our results regarding power and runtime. The results show that the proposed IMQ N-KSD approach has the highest power of all approximations across all tested $n$. Second best is L1 IMQ. W.r.t. runtime, the proposed method is faster than L1 IMQ for $n\in\{500,1000\}$. For $n=2000$, IMQ N-KSD has a similar runtime but still features better power. The FSSD approaches are the fastest but do not have a high power in this experiment.
Student-t vs. standard normal.
Again, we fix $d=15$, set $m=4\sqrt{n}$, and vary $n\in\{500,1000,2000\}$. The other parameters are the sames as the ones stated in Section 5.
Figure 2(b) shows that L1 IMQ achieves higher power than the proposed IMQ N-KSD for $n\in\{500,1000\}$ but at the price of a larger runtime. For $n=2000$, the performance of IMQ N-KSD is slightly better than that of L1 IMQ while both approaches have a similar runtime. The remaining approaches perform worse in terms of power.
Restricted Boltzmann machine (RBM).
For the RBM experiment, we set $\sigma=0.02$, $m=4\sqrt{n}$, and select $n\in\{500,1000,2000\}$; all other parameters match the ones detailed in Section 5.
We summarize the results in Figure 2(c). While both random feature Stein discrepancies (L1 IMQ, L2 SechExp) scale linearly in $n$, the higher dimensionality and difficulty of this problem result in a runtime that is orders of magnitude larger than that of all other approximations; the same observation w.r.t. runtime applies to the RFF approaches. We also observe that the runtimes of the related FSSD approaches increase compared to their runtime results in the Laplace and Student-t experiments.
Regarding power, the proposed Gauss N-KSD achieves the best result of all approximations from $n\geq 1000$ while being among the fastest methods. While, for $n\in\{1000,2000\}$, it is a bit slower than the FSSD approaches, the proposed method achieves higher power across all choices of $n$.
Summary. Figure 2(a)–(b) shows that some existing methods, e.g., L1 IMQ, perform similarly to N-KSD in terms of power achieved but come with a larger runtime for smaller sample sizes. Figure 2(c) highlights that some competitors (L1 IMQ, L2 SechExp, Cauchy RFF, Gauss RFF) require a larger runtime than the baseline approaches (IMQ KSD, Gauss KSD) for samples of size less than 2000. Here, our method is one to two orders of magnitude faster while achieving the same or, in some cases, larger power.
These results show that the proposed N-KSD has a very good runtime/power trade-off.
D.2 Impact of the Size of the Nyström Sample
Figure 3(a–d) captures the impact of the choice of Nyström samples $m=c\sqrt{n}$ for $c\in\{1,4,8\}$; the $\sqrt{n}$ dependence follows from Corollary 1(ii), where we neglect the logarithmic terms due to their small contribution. We include the quadratic time approaches as baselines; the experimental setup matches the experiments detailed in the article in Section 5. Generally, as one expects, both runtime and power increase for larger $c$. Still, even for $c=8$, where the power of the proposed approximation is hardly discernible from the baselines across all experiments, its runtime is an order of magnitude lower, which further strengthens the benefit of employing our proposed method.
Figure 3: Impact of different choices of factor $c$ for the number of Nyström samples $m=c\sqrt{n}$.